OpenStudy (anonymous):

Use the limit definition to find the derivitive of f(x)=1/(square root of x+1)

OpenStudy (anonymous):

\[f(x) = \frac{1}{\sqrt{x+1}}\]

OpenStudy (anonymous):

lim as x aproaches 0 of \[1\div \sqrt{x+1}\]

OpenStudy (anonymous):

Is the limit Definition like this \[\lim_{h \to 0} \:\:\:\:\:f(x+h) -f(x)\:\:?\]

OpenStudy (anonymous):

I am confused about it, If you can tell me a little about it then I might be able to help you.

OpenStudy (anonymous):

Use the limit definition above to solve for the derivative of f(x)

OpenStudy (anonymous):

\[f(x +h ) = \frac{1}{\sqrt{x +h + 1 } }\] \[\lim_{h \to 0 } \frac{1}{\sqrt{x+h+1}} - \frac{1}{\sqrt{x+1}}\] \[\lim_{h \to 0 } \frac{\sqrt{x+1} - \sqrt{x+h+1}}{\sqrt{x+h+1}\times\sqrt{x+1}} \]

OpenStudy (anonymous):

Now Rationalization of Numerator...

OpenStudy (anonymous):

\[\lim_{h \to 0 } \frac{x +1 - x -h -1 }{ \sqrt{x+h+1}\times{\sqrt{x+1}(\sqrt{x+1} + {\sqrt{x+h+1}})}} \]

OpenStudy (anonymous):

definition of the derivative is all over h noob

OpenStudy (anonymous):

\[\lim_{h \to 0 } \frac{-h }{ \sqrt{x+h+1}\times{\sqrt{x+1}(\sqrt{x+1} + {\sqrt{x+h+1}})}}\]

OpenStudy (anonymous):

should it all be over h so the h's cancel out?

OpenStudy (anonymous):

oh sorry I told you I am confused about it sorry yea h cancels out

OpenStudy (anonymous):

boob

OpenStudy (anonymous):

did I do something wrong now?

OpenStudy (anonymous):

no, i think youre doing everything right, dont listen to that "The_mathster"

OpenStudy (anonymous):

\[\lim_{h \to 0 } \frac{-1 }{ \sqrt{x+h+1}\times{\sqrt{x+1}(\sqrt{x+1} + {\sqrt{x+h+1}})}}\]h cancels out...And we have.. \[-\frac{1}{x +1 (2\sqrt{x+1})}\] \[-\frac{1}{2}\times \frac{1}{(x+1)^{\frac{3}{2}}}\]

OpenStudy (anonymous):

Limit complicates it, Differentiating is much easier...

OpenStudy (anonymous):

yah it is, but our assignment asks for us to use limit definition. thanks for your help